Can someone show me how to solve 64x^4 - 36x^2 = 0
Answer by scott8148(6628)

factoring gives (x^2)(64x^2-36)=0 ... factoring again gives (x)(x)(8x+6)(8x-6)=0

x=0 ...
8x+6=0 ... 8x=-6 ... x=-6/8 or -3/4 ...
8x-6=0 ... 8x=6 ... x=6/8 or 3/4